# Competing charge density wave phases in YNiC2

### v1

@@ Abstract
Charge density wave (CDW) orders in YNiC2 are studied by means of combined experimental and computational techniques. On the experimental side, single crystals grown by the floating-zone method were examined by means of X-ray diffraction, as well as transport and thermal techniques. Density functional theory (DFT) calculations founded on the experimentally determined parent and CDW-modified crystal structures provide details of electronic and phononic structures as well as electron-phonon coupling and resolve changes inflicted upon entering the different CDW phases. Thereby, contrasting effects of subsequently emerging CDW states characterized by incommensurate $q_{1ic}$ and commensurate $q_{2c}$ modulation vectors are revealed. The former state, on-setting below $T_{\text{1ic}}\simeq 305$ K, weakly modifies the electronic structure by opening an almost isotropic gap on a minor part of the Fermi surface (FS). The latter phase, which takes over below $T_{\text{2c}}\simeq 272$ K has a more pronounced impact on physical properties via a decomposition of larger parts of the FS. These dissimilar behaviors are directly reflected in the electronic transport anisotropy, which is significantly weakened in the $q_{2c}$-type CDW state. As revealed by our DFT studies, CDW phases are very close in energy and their origin is directly related to the anisotropy of electron-phonon coupling, which is linked to a specific orbital character of related FS sheets. Specific heat and thermal expansion studies reveal a nearly reversible first-order phase transition at around $T_{\text{2c}}\simeq 272$ K, where both CDW phases co-exist within a $T$-interval of about 10 K.
@@ I Introduction
An abundant sandbox for exploring interactions between various types of order parameters is provided by rare-earth nickel dicarbides $R$NiC2 – a class of materials known for a rich phase diagram comprising CDW phases, various magnetic states and superconductivity, systematically evolving with the lanthanide contraction [20, 21, 22, 23, 24, 25, 26]. Their electronic structure with quasi-one-dimensional character creates favorable conditions for CDW formation, in some cases with multiple CDW transitions. In $R$NiC2 compounds ($R$ = Pr – Sm) with larger unit-cell volume, CDW superstructures are characterized by a single incommensurate $q_{1ic}=(0.5,0.5+\delta,0)$ modulation wave vector, which may finally lock-in to a commensurate modulation $q_{1c}=(0.5,0.5,0)$ at low temperature as reported for GdNiC2 [20, 21, 22]. Upon replacing the rare-earth with heavier $R$-elements, one observes the emergence and eventually a prevalence of CDW order characterized by a commensurate modulation vector ${q_{2c}}=(0.5,0.5,0.5)$ [24, 25, 26]. YNiC2 hosts both, the incommensurate $q_{1ic}$-type and the commensurate $q_{2c}$-type CDW states [24]. Despite the observed signatures, the driving forces governing these two electronic instabilities have not been identified and explored yet.
@@ II Experimental
The polycrystalline material, needed to grow a single crystal, has been synthesized using radio-frequency induction technique in a high-purity argon atmosphere (99.9999%). Pure elements: Y (99.9%), Ni (99.99%) metals, and carbon (graphite, 99.999%) have been used as precursors in this procedure. In a next step, polycrystalline feed and seed rods were prepared, using the same induction melting method, and then used to grow monocrystalline YNiC2 via the floating zone technique in an optical mirror furnace (Crystal Systems Corporation, Japan). Single crystals were finally oriented by means of the Laue method and cut along the principal orientations of the orthorhombic parent structure, in dimensions as desired for specific measurements.
@@
Table 1: Data collection and refinement details of the incommensurately modulated and triperiodic CDW phases of YNiC2. $R$- and $wR2$-values of the incommensurate structure at 280 K refer to refinements of all reflections, main reflections and $1^{\rm st}$ order satellites, respectively.
@@
   measured 2780 2780 8226
   unique main 378 467 1659
    $1^{\rm st}$ order satellites 794
   observed ($I>2\sigma(I)$) 377 467 960
    $1^{\rm st}$ order satellites 469
@@
The special points for band structures and phonon dispersions for the parent phase are defined according to Ref. [35]: $\Gamma=(0.00,0.00,0.00)$; $S=(0.00,0.50,0.00)$; $R=(0.00,0.50,0.50)$; $Z=(0.00,0.00,0.50)$; $T=(-0.50,0.50,0.50)$; $Y=(-0.50,0.50,0.00)$; $\Sigma_{0}=(0.39,0.39,0.00)$; $A_{0}=(0.39,0.39,0.50)$ (see Fig. 7a). The details on how we obtained the corresponding points for the $q_{1c}$- and $q_{2c}$-type CDW cells are given in SM [34] Sect. II.
@@ III.1 Crystal structures
Like all members of the $R$NiC2 family at elevated temperatures, YNiC2 also crystallize in the orthorhombic $Amm2$ aristotype structure (Fig. 1a) [44]. All metal atoms and the C2 dumbbells are located on sites with $mm2$ symmetry. The Ni atoms and the C2 dumbbells are located on the $x=0$ planes, the Y atoms on the $x=\frac{1}{2}$ planes. The Ni and Y atoms form linear chains extending along $[100]$, whereby adjacent atoms are related by the $\mathbf{a}$ lattice translation with interatomic distances at 320 K, $a=3.5724$ (2) Å.
@@
Between $T_{\textbf{1ic}}$ and $T_{\mathbf{2c}}$, YNiC2 adopts a 1D (along the orthorhombic $b$-axis) incommensurately modulated structure analogous to SmNiC2 [45] with superspace group symmetry $Amm2(\frac{1}{2}\sigma_{2}0)000$, and modulation wave vector $\mathbf{q}_{1ic}=\frac{1}{2}\mathbf{a}^{*}+\sigma_{2}\mathbf{b}^{*}$ with $\sigma_{2}=0.5138(8)$ at 280 K. Accordingly, periodicity in the $[100]$ direction is halved (doubling of the $a$-axis, as in the monoclinic $Cm$ structures) fully retained in the $[001]$ direction and lost in $[010]$ direction. For single crystal XRD data taken right at 275 K, i.e. in the vicinity of $T_{2c}$, satellite reflections related to both modulation wave vectors $q_{2c}$ and $q_{1ic}$ are observed and attributed to a coexistence of $q_{2c}$- and $q_{1ic}$-type CDW domains (see Sect. III.3 for a discussion on details of the related phase transition and Fig. S1 in SM [34], for the diffraction image).
Since the basic structures of the incommensurate phases are isotypic to the aristotype phases, comparable atomic coordinates were used, up to the $x$-coordinates, which where shifted by $\frac{1}{2}$. The $Amm2$ basic structure contains two symmetrically equivalent reflection planes parallel to $(100)$ (see Fig. 1a) and the origin may be chosen on either of the planes. In the $Amm2(\frac{1}{2}\sigma_{2}0)000$ superspace group, these reflection planes split in two distinct symmetry elements. The $x=0$ plane remains a proper mirror plane in the modulated structure. In contrast, reflecting at $x=\frac{1}{2}$ shifts the modulation functions by half a period, which can be regarded as the incommensurate equivalent of a glide reflection. If the $x$-coordinates of the basic structure were retained (Ni at $x=0$), the Ni atoms could not be displaced in the $[100]$ direction. The resulting model gives refinements with excellent reliability factors, but is unlikely, given the Peierls pair formation in the commensurate $Cm$ structures. Note that an analogous phenomenon is observed in the commensurate LT phase: The reflection planes parallel to $(100)$ of the aristotype structure split into an alternation of proper reflection planes and glide reflection planes with an intrinsic translation component of $\frac{1}{2}(\mathbf{b}+\mathbf{c})$ [25]. There as well, the Ni atoms had to be moved to $x=\frac{1}{2}$.
@@
By subtraction of Eqns. 1 and III.1, one notes that in the $[100]$ direction, Ni atoms are arranged in lines spaced alternately by
@@
Figure 3: (a) Temperature-dependent electrical resistivity, $\rho(T)$, measured on bar-shape YNiC2 single crystals which were cut along orthorhombic $a$-, $b$- and $c$-orientations; (b) normalized electrical resistivity $\rho(T)/\rho(T_{1ic})$; (c) temperature-dependent relative anisotropy of the electrical resistivity ${\rho_{b}}/{\rho_{a}}$ and ${\rho_{b}}/{\rho_{a}}$.
@@ III.2 Electrical resistivity studies
A charge density wave is essentially a coupled lattice and electronic instability, typically associated with the opening of one or more gaps at the Fermi energy. The evolution of the electronic component is directly reflected in the charge transport properties, which can be explored by macroscopic observables, such as the temperature-dependent electrical resistivity, $\rho(T)$. The latter, i.e. $\rho_{a}$, $\rho_{b}$, $\rho_{c}$, measured for single crystalline YNiC2 with applied current along principal orientations is displayed in Fig. 3a and reveals distinct anomalies related to the two CDW transitions discussed in Sect. III.1. The first transition at $T_{1ic}=305$ K (i) from the orthorhombic CeNiC2-type parent structure to the orthorhombic incommensurate $q_{1ic}$-type CDW and the second transition at $T_{2c}=272$ K (ii) into a monoclinic commensurate $q_{2c}$-type CDW phase are well reflected by characteristic anomalies of the electrical resistivity, kink-like at $T_{1ic}$ and jump-like at $T_{2c}$, respectively.
At temperatures above $T_{1ic}$, $\rho(T)$ data of YNiC2 in Fig. 3a share the features of an anisotropic electrical resistivity $\rho_{c}>\rho_{b}>\rho_{a}$ as earlier reported for other $R$NiC2 compounds in their CeNiC2-type parent state [21, 22, 25, 26]. At transition (i), a clear minimum at $T_{1ic}$ followed by a $\rho(T)$ increase upon lowering the temperature, is observed for all three orthorhombic orientations (compare normalized resistivity data in Fig. 3b and relative anisotropy data in Fig. 3c), thus, revealing an almost unchanged anisotropy of $\rho_{a}$ vs. $\rho_{b}$ vs. $\rho_{c}$ across $T_{1ic}$. The latter is a characteristic feature of the second-order phase transition into the $q_{1ic}$-type CDW state as previously reported for other $R$NiC2 (R = Sm, Gd, Tb) representatives exhibiting this specific $q_{1ic}$-type CDW transition with a preserved orthorhombic point group symmetry [21, 22].
@@
The transition (ii) from the $q_{1ic}$-type into the commensurate $q_{2c}$-type CDW state at $T_{2c}$ is manifested by jump-like anomalies of all three $\rho(T)$ curves in Fig. 3a, i.e. by visible inflections seen simultaneously, yet clearly different for the different orientations. While $\rho_{b}(T)$ and $\rho_{c}(T)$ show a sharp downturn, the resistivity measured along the $a$ direction, previously weakly affected at $T_{1ic}$, now visibly increases upon lowering the temperature. The pronounced anomaly of $\rho_{a}(T)$ accompanying the CDW crossover into the $q_{2c}$-type CDW state is not a sole domain of YNiC2, but appears to be a characteristic feature for this type of CDW order. Similarly specific response of $\rho_{a}(T)$ has been previously observed for single-crystalline TmNiC2 and LuNiC2 with a single CDW transition directly to the $q_{2c}$-type CDW state [25, 26]. With further decrease of temperature, the electrical resistivity gradually decreases for all three orientations. Within 100 – 200 K, a visible change of electronic anisotropy takes place with both ${\rho_{b}}/{\rho_{a}}$ and ${\rho_{b}}/{\rho_{a}}$ ratios gradually decreasing and saturating closer to unity as T is lowered (see Fig. 3c). A rather weak anisotropy of the electrical resistivity remains at lowest temperatures, with the lowest value for $\rho_{b}$, similar to earlier reported data of TmNiC2 and LuNiC2 [25, 26]. The residual resistivity of YNiC2, as measured at 2 K, is $\rho_{a0}$ = 27 $\mu\Omega$cm, $\rho_{b0}$ = 20 $\mu\Omega$cm, $\rho_{c0}$ = 33 $\mu\Omega$cm.
Motivated by a previous transport study of polycrystalline YNiC2, which revealed the presence of a finite thermal hysteresis opening approximately at 285 K and closing at 265 K [39], we have investigated the character of both CDW anomalies in our single crystal via an electrical resistivity measurement with specific care on an optimal temperature stabilization (see SM [34], Sect. III for further details). The corresponding $\rho_{c}(T)$ measurement (see Fig. S2 in SM [34]) clearly confirms thermal reversibility of the electrical resistivity upon heating and cooling across the second order transition (i) at $T_{1ic}$ and reveals very narrow, but nonetheless finite thermal hysteresis of about 0.1 K for the anomaly at around $T_{2c}$. From the double-peak-like character of the temperature derivative of $\rho(T)$, we determine the width of transition (ii) ranging from about 265 to 275 K (see Fig. S3a in SM [34]). Accordingly, the width of the transition ($\sim 10$ K) reveals to be two orders of magnitude larger than the width of the thermal hysteresis ($\sim 0.1$ K) and the transition is thus almost perfectly reversible.

### v2

@@ Abstract
Charge density wave (CDW) orders in YNiC2 are studied by means of combined experimental and computational techniques. On the experimental side, single crystals grown by the floating-zone method were examined by means of X-ray diffraction, as well as transport and thermal techniques. Density functional theory (DFT) calculations founded on the experimentally determined parent and CDW-modified crystal structures provide details of electronic and phononic structures as well as electron-phonon coupling, and resolve changes inflicted upon entering the different CDW phases. Thereby, contrasting effects of subsequently emerging CDW states characterized by incommensurate $q_{1ic}$ and commensurate $q_{2c}$ modulation vectors are revealed. The former state, on-setting below $T_{\text{1ic}}\simeq 305$ K, weakly modifies the electronic structure by opening an almost isotropic gap on a minor part of the Fermi surface (FS). The latter phase, which takes over below $T_{\text{2c}}\simeq 272$ K has a more pronounced impact on physical properties via a decomposition of larger parts of the FS. These dissimilar behaviors are directly reflected in the electronic transport anisotropy, which is significantly weakened in the $q_{2c}$-type CDW state. As revealed by our DFT studies, CDW phases are very close in energy and their origin is directly related to the anisotropy of electron-phonon coupling, which is linked to a specific orbital character of related FS sheets. Specific heat and thermal expansion studies reveal a nearly reversible first-order phase transition at around $T_{\text{2c}}\simeq 272$ K, where both CDW phases co-exist within a $T$-interval of about 10 K.
@@ I Introduction
An abundant sandbox for exploring interactions between various types of order parameters is provided by rare-earth nickel dicarbides $R$NiC2 – a class of materials known for a rich phase diagram comprising CDW phases, various magnetic states and superconductivity, systematically evolving with the lanthanide contraction [20, 21, 22, 23, 24, 25, 26]. Their electronic structure with quasi-one-dimensional character creates favorable conditions for CDW formation, in some cases with multiple transitions. In $R$NiC2 compounds ($R$ = Pr – Sm) with larger unit-cell volume, CDW superstructures are characterized by a single incommensurate $q_{1ic}=(0.5,0.5+\eta,0)$ modulation wave vector, which may finally lock-in to a commensurate modulation $q_{1c}=(0.5,0.5,0)$ at low temperature as reported for GdNiC2 [20, 21, 22]. Upon replacing the rare-earth with heavier $R$-elements, one observes the emergence and eventually a prevalence of CDW order characterized by a commensurate modulation vector ${q_{2c}}=(0.5,0.5,0.5)$ [24, 25, 26]. YNiC2 hosts both, the incommensurate $q_{1ic}$-type and the commensurate $q_{2c}$-type CDW states [24]. Despite the observed signatures, the driving forces governing these two electronic instabilities have not been identified and explored yet.
@@ II Experimental
The polycrystalline material, needed to grow a single crystal, has been synthesized using radio-frequency induction melting in a high-purity argon atmosphere (99.9999%). Pure elements: Y (99.9%), Ni (99.99%) metals, and carbon (graphite, 99.999%) have been used as precursors in this procedure. In a next step, polycrystalline feed and seed rods were prepared, using the same induction melting method, and then used to grow monocrystalline YNiC2 via the floating zone technique in an optical mirror furnace (Crystal Systems Corporation, Japan). Single crystals were finally oriented by means of the Laue method and cut along the principal orientations of the orthorhombic parent structure, in dimensions as desired for specific measurements.
@@
Table 1: Data collection and refinement details of the incommensurately modulated and triperiodic CDW phases of YNiC2. $R$- and $wR2$-values of the modulated structures at 280 and 250 K refer to refinements of all reflections, main reflections and $1^{\rm st}$ order satellites, respectively. For the commensurate structure at 250 K, the main reflections are defined as those with even $k$. $R_{\mathrm{int}}$ is given for all reflections (main and satellites).
@@
   measured 2780 4548 8226
   unique main 378 467 583
    $1^{\rm st}$ order satellites 794 1073
   observed ($I>2\sigma(I)$) 377 467 543
    $1^{\rm st}$ order satellites 469 414
@@
The special points for band structures and phonon dispersions for the parent phase are defined according to Ref. [35]: $\Gamma=(0.00,0.00,0.00)$; $S=(0.00,0.50,0.00)$; $R=(0.00,0.50,0.50)$; $Z=(0.00,0.00,0.50)$; $T=(-0.50,0.50,0.50)$; $Y=(-0.50,0.50,0.00)$; $\Sigma_{0}=(0.39,0.39,0.00)$; $A_{0}=(0.39,0.39,0.50)$. The details on how we obtained the corresponding points for the $q_{1c}$- and $q_{2c}$-type CDW cells are given in SM [34] Sect. II.
@@ III.1 Crystal structures
Like all members of the $R$NiC2 family at elevated temperatures, YNiC2 also crystallizes in the orthorhombic $Amm2$ aristotype structure (Fig. 1a) [44]. All metal atoms and the C2 dumbbells are located on sites with $mm2$ symmetry. The Ni atoms and the C2 dumbbells are located on the $x=0$ planes, the Y atoms on the $x=\frac{1}{2}$ planes. The Ni and Y atoms form linear chains extending along $[100]$, whereby adjacent atoms are related by the $\mathbf{a}$ lattice translation with interatomic distances at 320 K, $a=3.5724$ (2) Å.
@@
Between $T_{\textbf{1ic}}$ and $T_{\mathbf{2c}}$, YNiC2 adopts a 1D (along the orthorhombic $b$-axis) incommensurately modulated structure analogous to SmNiC2 [45] with superspace group symmetry $Amm2(\frac{1}{2}\sigma_{2}0)000$, and modulation wave vector $\mathbf{q}_{1ic}=\frac{1}{2}\mathbf{a}^{*}+\sigma_{2}\mathbf{b}^{*}$ with $\sigma_{2}=0.5138(8)$ at 280 K. Accordingly, periodicity in the $[100]$ direction is halved (doubling of the $a$-axis, as in the monoclinic $Cm$ structures) fully retained in the $[001]$ direction and lost in $[010]$ direction. For single crystal XRD data taken right at 275 K, i.e. in the vicinity of $T_{2c}$, satellite reflections related to both modulation wave vectors $q_{2c}$ and $q_{1ic}$ are observed and attributed to a coexistence of $q_{2c}$- and $q_{1ic}$-type CDW domains (see Sect. III.3 for a discussion on details of the related phase transition and Fig. S1 in SM [34], for the diffraction image). Observe that the $q_{2c}$- and $q_{1ic}$ modulation wave vectors possess fundamentally different $\mathbf{c}$ components ($\frac{1}{2}$ and 0, respectively) and therefore the structures cannot be related by a unified superspace description. Thus, technically, the $q_{1ic}$ to $q_{2c}$ transition is not of lock-in type.
Since the basic structure of the incommensurate phase is isotypic to the aristotype phase, comparable atomic coordinates were used, up to the $x$-coordinates, which were shifted by $\frac{1}{2}$. The $Amm2$ basic structure contains two symmetrically equivalent reflection planes parallel to $(100)$ (see Fig. 1a) and the origin may be chosen on either of the planes. In the $Amm2(\frac{1}{2}\sigma_{2}0)000$ superspace group, these reflection planes split in two distinct symmetry elements. The $x=0$ plane remains a proper mirror plane in the modulated structure. In contrast, reflecting at $x=\frac{1}{2}$ shifts the modulation functions by half a period, which can be regarded as the incommensurate equivalent of a glide reflection. If the $x$-coordinates of the basic structure were retained (Ni at $x=0$), the Ni atoms could not be displaced in the $[100]$ direction. The resulting model gives refinements with excellent reliability factors, but is unlikely, given the Peierls pair formation in the commensurate $Cm$ structures. Note that an analogous phenomenon is observed in the commensurate LT phase: The reflection planes parallel to $(100)$ of the aristotype structure split into an alternation of proper reflection planes and glide reflection planes with an intrinsic translation component of $\frac{1}{2}(\mathbf{b}+\mathbf{c})$ [25]. There as well, the Ni atoms had to be moved to $x=\frac{1}{2}$.
@@
By subtraction of Eqns. 1 and III.1, one notes that in the $[100]$ direction Ni atoms are arranged in lines spaced alternately by
@@
Figure 3: (a) Temperature-dependent electrical resistivity, $\rho(T)$, measured on bar-shape YNiC2 single crystals which were cut along orthorhombic $a$-, $b$- and $c$-orientations; (b) normalized electrical resistivity $\rho(T)/\rho(T_{1ic})$; (c) temperature-dependent relative anisotropy of the electrical resistivity ${\rho_{b}}/{\rho_{a}}$ and ${\rho_{c}}/{\rho_{a}}$.
@@ III.2 Electrical resistivity studies
A charge density wave is essentially a coupled lattice and electronic instability, typically associated with the opening of one or more gaps at the Fermi energy. The evolution of the electronic component is directly reflected in the charge transport properties, which can be explored by macroscopic observables, such as the temperature-dependent electrical resistivity, $\rho(T)$. The latter, i.e. $\rho_{a}$, $\rho_{b}$, $\rho_{c}$, measured for single crystalline YNiC2 with current applied along principal orientations is displayed in Fig. 3a and reveals distinct anomalies related to the two CDW transitions discussed in Sect. III.1. The first transition at $T_{1ic}=305$ K (i) from the orthorhombic CeNiC2-type parent structure to the orthorhombic incommensurate $q_{1ic}$-type CDW and the second transition at $T_{2c}=272$ K (ii) into a monoclinic commensurate $q_{2c}$-type CDW phase are well reflected by characteristic anomalies of the electrical resistivity, kink-like at $T_{1ic}$ and jump-like at $T_{2c}$, respectively.
At temperatures above $T_{1ic}$, $\rho(T)$ data of YNiC2 in Fig. 3a share the features of an anisotropic electrical resistivity $\rho_{c}>\rho_{b}>\rho_{a}$ as earlier reported for other $R$NiC2 compounds in their CeNiC2-type parent state [21, 22, 25, 26]. At transition (i), a clear minimum at $T_{1ic}$, followed by a $\rho(T)$ increase upon lowering the temperature, is observed for all three orthorhombic orientations (compare normalized resistivity data in Fig. 3b and relative anisotropy data in Fig. 3c), thus, revealing an almost unchanged anisotropy of $\rho_{a}$ vs. $\rho_{b}$ vs. $\rho_{c}$ across $T_{1ic}$. The latter is a characteristic feature of the second-order phase transition into the $q_{1ic}$-type CDW state as previously reported for other $R$NiC2 ($R=$ Sm, Gd, Tb) representatives exhibiting this specific $q_{1ic}$-type CDW transition with a preserved orthorhombic point group symmetry [21, 22].
@@
The transition (ii) from the $q_{1ic}$-type into the commensurate $q_{2c}$-type CDW state at $T_{2c}$ is manifested by jump-like anomalies of all three $\rho(T)$ curves in Fig. 3a, i.e. by visible inflections seen simultaneously, yet clearly different for the different orientations. While $\rho_{b}(T)$ and $\rho_{c}(T)$ show a sharp downturn, the resistivity measured along the $a$ direction, previously weakly affected at $T_{1ic}$, now visibly increases upon lowering the temperature. The pronounced anomaly of $\rho_{a}(T)$ accompanying the CDW crossover into the $q_{2c}$-type CDW state is not a sole domain of YNiC2, but appears to be a characteristic feature for this type of CDW order. Similarly specific response of $\rho_{a}(T)$ has been previously observed for single-crystalline TmNiC2 and LuNiC2 with a single CDW transition directly to the $q_{2c}$-type CDW state [25, 26]. With further decrease of temperature, the electrical resistivity gradually decreases for all three orientations. Within 100 – 200 K, a visible change of electronic anisotropy takes place with both ${\rho_{b}}/{\rho_{a}}$ and ${\rho_{c}}/{\rho_{a}}$ ratios gradually decreasing and saturating closer to unity as T is lowered (see Fig. 3c). A rather weak anisotropy of the electrical resistivity remains at lowest temperatures, with the lowest value for $\rho_{b}$, similar to earlier reported data of TmNiC2 and LuNiC2 [25, 26]. The residual resistivity of YNiC2, as measured at 2 K, is $\rho_{a0}$ = 27 $\mu\Omega$cm, $\rho_{b0}$ = 20 $\mu\Omega$cm, $\rho_{c0}$ = 33 $\mu\Omega$cm.
Motivated by a previous transport study of polycrystalline YNiC2, which revealed the presence of a finite thermal hysteresis opening approximately at 285 K and closing at 265 K [39], we have investigated the character of both CDW anomalies in our single crystal via an electrical resistivity measurement with specific care on an optimal temperature stabilization (see SM [34], Sect. III for further details). The corresponding $\rho_{c}(T)$ measurement (see Fig. S2 in SM [34]) clearly confirms thermal reversibility of the electrical resistivity upon heating and cooling across the second order transition (i) at $T_{1ic}$ and reveals for transition (ii) a very narrow, but nonetheless finite thermal hysteresis of about 0.1 K for the anomaly at around $T_{2c}$. From the double-peak-like character of the temperature derivative of $\rho(T)$, we determine the width of transition (ii) ranging from about 265 to 275 K (see Fig. S3a in SM [34]). Accordingly, the width of the transition ($\sim 10$ K) reveals to be two orders of magnitude larger than the width of the thermal hysteresis ($\sim 0.1$ K) and the transition is thus almost perfectly reversible.